Cod liver oil is available in capsules containing 0.6 mL per capsule. Vitamin A : 1USP Vitamin A unit equals the biologic activity of 0.3mcg of the all-trans isomer of retinol and Vitamin D - 40 units per mcg, calculate the amounts, in units, each of vitamins A and D in each capsule. The specific gravity of cod liver oil is 0.92.
Answer #1

The volume of one capsule is 0.6 mL, thus the calculations can be made as below:

Amount of vitamin D in the capsule = density * volume, or 0.92 * 1000 * 0.6 = 0.552 g or 552 mg

1 microgram or 10-6 g of vitamin D = 1 USP

thus, units of vitamin D in one capsule = 552 / 10-6 = 5.52 * 105 g

One unit of vitamin D has biological acitivyt = 0.3 microgram

Thus, the total units ofbiological activities obtained = 0.3 * 5.52 * 105 or 1.656 *105 mcg.